A metallic sphere floats in immiscible mixture of water (density 103 kg/m3) and a liquid(density 8X103 kg/m3) such that its (2/3) part is in water (1/3) part in the liquid.The density of the metal is


a.5000/3 kg/m3


b.10000/3 kg/m3


c.5000 kg/m3


d.2000 kg/m3

Grade:12th Pass

1 Answers

nitish kumar
12 years ago

for floation weight of the sphere should be equal to the net bouyant force

d=density of sphere ,d1=density of water,d2 =density of liquid

bouyant force f= weightof water displaced+weight of liquid displaced

f={(2/3)V*d1*g]+[(1/3)V*d2*g] = V*d*g

solving this we will get d=10000/3 kg/m3